if the selling price per unit is $70, the variable expense per unit is $50, and total fixed expenses are $261,000, what are the breakeven sales in dollars?     

I'm getting 913,545 but that's not an option
Answer #1

Contribution margin ratio = (70-50)/70 = 28.5714%

Breakeven in sales dollars = Fixed expenses / Contribution ratio

= 261000 / 0.285714 = $913,500

I'm getting 913,545 but that's not an option - May be because or rounding of you are getting this.